### v2

Abstract
[…]
Main
The field-of-view (FOV) of an optical system denotes its maximum observable area, which, together with resolution, is crucial in live-cell imaging for monitoring cell and tissue dynamics for biomedical applications such as drug screening and disease research. Most of the current biomedicine imaging modalities rely on light microscopy, wherein light is either transmitted or reflected by the sample through lenses. However, due to the lens’ magnification by the numerical aperture (NA) and the sensor size, FOV in lens-based optical systems is inherently limited [1]. Efforts to expand the FOV have historically focused on advancements in lens design and reconstruction algorithms [2, 3, 4, 5]. For lens-based microscopes, despite significant enhancements in contrast [6], specificity [7, 8], optical sectioning capabilities [9, 10], and advances in super-resolution algorithms [11, 12, 13, 14, 15], FOV remains constrained by magnification and lens aberrations, with commercial systems offering a maximum FOV of $312.5\,mm^{2}$ at minimal resolution[16]. Consequently, existing biomedical imaging platforms, like microscopes and plate readers, require scanning for a larger FOV by regularly adjusting lens positions. This intrinsic limitation impedes real-time imaging on large tissue sections (i.e. metastatic cancer processes across microvessels), whole model organisms (e.g., zebrafish), and advanced cell models (e.g., spheroids/organoids and organs on chip).
To overcome these limitations, lens-free imaging (LFI) strategies have been developed. LFI, a form of computational imaging (CI), captures diffraction patterns (i.e., holograms) directly with the image sensor and reconstructs focused images using computational algorithms, bypassing the need for conventional lenses. Like other CI systems, current LFI reconstruction algorithms depend on a physical forward model for a resolvable optical field. In 2D LFI systems [17, 18, 19, 20, 21, 22, 23], the distance between the sample and the light source plays a vital role in inverse computation for the diffraction field. This parameter limits the imaging of the sample to a thin focal plane. In parallel, 3D tomographic LFI systems [24, 25, 26], which rely on multi-position measurements, offer a degree of depth of field in the range of micrometers. Nevertheless, these algorithms inherently demand a high-quality and resolvable hologram, which, in real-world scenarios, can only come from a clean, non-overlapping, and stable diffraction optical field modeling. This necessitates precise control over the samples’ size and density to align with precise propagation parameters, such as depth of field and illumination angle [27]. Although deep-learning (DL) has improved holographic reconstruction (re-focusing) algorithms enabling their generalization to diverse sample types [17, 28, 29], such idealized measurement conditions limit the applicability of current LFI systems to imaging of static, specific and well-controlled samples. Moreover, the sensitivity of their optical instruments to environmental factors such as vibration, temperature, and contamination further limits their practical use in dynamic, real-world scenarios [24, 20, 30]. Finally, to achieve a FOV and resolution similar to confocal microscopes, current LFI systems [27, 30] usually require a shifting light source to scan a large region-of-interest (ROI). Because of all these limitations, state-of-the-art LFI systems currently offer little characterization advantages over traditional optical imaging techniques.
Existing biomedical imaging systems all encounter a significant constraint: limited throughput through one-shot FOV. However, a modest FDA-approved drug library can contain around 2000 samples and 2 million or more pairwise combinations [31], where the time delay to image different samples heavily decreases the ability to interrogate dynamic events. Notably, the repetitive nature of these samples’ high-content data presents an opportunity. In the context of high-throughput imaging for drug screening, previous studies [32, 33, 34, 35, 36] have demonstrated the reliability of deep neural networks, such as generative adversarial networks (GAN). By leveraging these recurrent patterns as robust data guidance, it is possible to enhance the reconstruction of sample morphology more effectively, thus mitigating FOV limitations and enhancing the efficiency of high-throughput imaging.
Here, we propose a generative deep learning-enabled LFI framework, namely GenLFI, employing DL-based computer vision techniques to address the limitations of state-of-the-art LFI systems. GenLFI employs a universal optical setup and significantly simplifies the imaging process by eliminating the need for data pre-processing, annotation, and optical field modeling that traditional LFI systems depend on. Our framework also incorporates an unsupervised learning model, LensGAN, capable of autonomously learning image features from unpaired and unaligned holograms and bright-field light microscopy to generate high-resolution image reconstructions. This is achieved through a comprehensive analysis of diffraction and projection principles, establishing physical correlations between holograms and reconstructed bright-field images.
Through a series of imaging experiments, we demonstrate how the GenLFI framework advances LFI systems in three key aspects: (1) Our LensGAN’s robust learning capability enables the reconstruction of high-quality images from low-resolution holograms, achieving an ultra-large FOV in a single shot. This FOV is solely constrained by the sensor size, with our validation on a full-frame CMOS sensor demonstrating a FOV exceeding $550\,mm^{2}$, while still preserving sub-sensor-pixel resolution. (2) Unlike state-of-the-art LFI methods, GenLFI requires no optical field control or modeling, allowing it to image complex 3D samples (i.e., microfluidics, 3D cell models). (3) LensGAN boasts a faster inference speed than current hologram reconstruction algorithms [17, 37, 38]. These advancements collectively extend the boundaries of LFI technology in high-throughput biomedical imaging tasks. The expanded, streamlined, and enhanced observational capacities of GenLFI can create new possibilities for broad dynamic imaging applications in complex cell-based microfluidic systems, phenotypic screening, and AI-enabled digital pathology.
Results
GenLFI framework
Figure 1: Overview of GenLFI framework. a, the left illustrates our ultra-large FOV LFI setup, and the right showcases examples of 3D samples. The middle depicts the complex optical field: as a coherent light source illuminates the sample, scattered waves interfere with the undistributed incident wave, forming a complex hologram above the CMOS sensor through the interference pattern and shaded projection. b, the workflow of GenLFI.
Incorporating physics-based priors is essential in LFI. Previous LFI reconstruction approaches rely on a forward imaging model, denoted by $H$, that relates the measured holograms $i$ to the reconstruction $\hat{o}$, by incorporating the physical principles of the LFI system [39]. The LFI hologram reconstruction task can thus be formulated as [17]:
$\hat{o}=arg\min_{o}L(H(o),i)+R(o),$ (1)
where $i\in\mathbb{H}^{IK^{2}}$ denotes the measured holograms after vectorization; $I$ represents the number of measured raw holograms, each with a dimension of $K\times K$; $o$ and $\hat{o}\in\mathbb{R}^{K^{2}}$ are the real-world ground truth and predicted target image with a dimension of $K\times K$ after vectorization, respectively, which in this study are bright-field images; $L(\cdot)$ and $R(\cdot)$ are the loss function and regularization term, respectively.
However, crafting an accurate physical forward model $H$ within a complex optical field, shown in Fig.1a, presents significant challenges due to several factors. The 3D sample, with vertical dimensions measured in micrometers ($\mu m$), can produce overlapped scattered waves from various axial distances and angles, particularly when interacting with mediums of different refractive indexes. In such scenarios, the optical field generates a complex hologram at the sensor, incorporating diffraction waves from different height planes and shadow projections. Furthermore, the optical field is subject to alterations caused by the movement of dynamic objects, which introduces additional distortions. While certain methods can enhance the robustness of light field modeling to perturbations in complex optical fields [36, 21], these factors collectively contribute to challenges in accurately describing, modeling, and reconstructing the optical field using current computational algorithms [22, 29, 18, 17, 19, 28, 24, 40] (further discussion see Supplementary Information 2).
To address these challenges, we introduce GenLFI, a generative deep learning-enabled LFI framework. It comprises a universal optical setup (see Methods), real-world data guidance, and a hologram reconstruction algorithm, i.e., LensGAN. As shown in Fig.1b, The workflow of GenLFI is as follows: (1) The unpaired dataset consists of the data acquired from our LFI setup (i.e., measured holograms) and the data from a bright-field microscope (i.e., real-world bright-field images), which are rescaled to serve as the input for LensGAN. (2) During training, LensGAN generalizes both the measured and reconstructed holograms ($i$ and $\hat{i}$) and the bright-field images ($o$ and $\hat{o}$) into two domains, $X$ and $Y$, respectively. (3) The trained LensGAN model is then deployed to rapidly reconstruct hologram patches, which are finally synthesized into original-sized holograms.
LensGAN adopts an unsupervised learning approach, bypassing the reliance on the forward model $H$ to decode complex optical fields. Instead, it leverages the power of generative models for robust image transformations within complex light fields. Simultaneously, LensGAN captures the inherent physical correlations between holograms and bright-field images. Our approach introduces two key enhancements to establish these physical prior correlations and optimize the model. First, we designed a discriminator network ($D$) that integrates one-layer attention between the downsampling convolutional layers to improve the processing of fine details in amplitude (e.g., illumination details). Second, we devise and incorporate a phase-contrast module to mimic the mechanism of the phase-contrast microscope (PCM) and dark-field microscope (DFM) due to the similar transparent nature of the objects (details in Methods). This module enforces physical consistency between the phase details (e.g., structures) of the hologram and the bright-field image. These improvements eliminate the need for optical field modeling required by existing LFI methods and manual annotation in typical supervised learning frameworks.
This mechanism liberates the optical setup from the constraints of optical field modeling, offering flexibility in setup design and robustness against environmental disturbance. We constructed a universal LFI setup adhering to the unit-magnification principle [41] (see Fig.1a). To capture an ultra-large FOV hologram in a single shot, we considerably extended the distance beyond that of conventional setups (details in meth-optset). We leverage this setup to demonstrate the flexibility, accuracy, efficiency, and robustness of the LensGAN algorithm in the following sections.
FOV, Resolution and Multi-focus imaging
FOV: In LFI, the achievable FOV captured in a single shot (one-shot FOV) is limited by the numerical aperture (NA) of the light source and the distance $f$ between the light source and the detector, e.g., the CMOS sensor in Fig.1a. While previous research[21, 28, 36] has demonstrated some flexibility in adjusting this distance within a millimeter ($mm$) range, achieving the best lateral resolution requires a specific optimized distance. This optimized distance [42] can be calculated as:
$f_{opt}=\frac{d\lambda}{Np-\frac{d\lambda^{2}}{Np}}$ (2)
where $\lambda$ is the parameter wavelength; $N$ is the number of pixels of the image sensor; $p$ denotes the pixel size, and $d$ is the distance between the object and sensor. The optimized distance is essential for generating high-resolution holograms, which is the cornerstone of existing LFI holographic reconstruction algorithms. Therefore, with a given NA, the FOV of existing LFI systems is constrained by the distance $f\approx f_{opt}$. LensGAN overcomes this limitation by reconstructing images directly from low-resolution holograms, eliminating the need for precise diffraction pattern decoding. This enables the GenLFI framework to surpass the $f_{opt}$ constraint, offering flexibility on the distance $f$. The FOV limitation is then primarily governed by the sensor size. In our demonstration, the GenLFI framework achieves a FOV of $553\,mm^{2}$ (see Fig.2a with the setup in Methods). Further improvements in FOV are possible through an upgraded optical configuration, such as using a higher NA light source, and a larger and higher-resolution image sensor.
Figure 2: FOV and resolution characterization. a shows different imaging systems’ FOV on our microfluidic chip design. 4x: 4x objective of bright-field microscope (details see Optical setup); state-of-the-art LFI [43]: the existing LFI system’s one-shot FOV based on holographic in-line reconstruction. AX 1x: Nikon AX [16], the world’s largest commercial confocal microscope. The scanning 1x objective can cross a $25\,mm$ diagonal FOV in multiple shots. b, Imaging of the USAF 1951 resolution test chart. The input hologram was paired with the ROI of the smallest region of the test chart in the bright field. The green box indicates the reconstructed resolvable region; the curves of the cross-sectional profile at the bottom row demonstrate the distinguishable line pairs.
Resolution: The image resolution attainable by GenLFI is affected by two physical parameters. First, the Nyquist-Shannon sampling theorem dictates that the lateral resolution of holograms is inherently constrained by the pixel size of the CMOS sensor, which is $5.94\,\mu m$ in our setup. This limitation suggests that details finer than the sensor’s pixel pitch cannot be adequately captured in the hologram [30]. Second, the reconstructed image quality is significantly influenced by the dimension of the input image of the model (see Methods). In particular, LensGAN’s training is memory-intensive on GPU due to the need to load six networks (see Methods), limiting the input image dimensions due to hardware constraints. We imaged a 1951 USAF resolution test chart and conducted a cross-sectional profile analysis following the standard routine [44, 45, 46]. We selected an area within the measured hologram where the smallest line bar is not resolvable (see Fig.2b). After reconstruction, we found that the smallest resolvable line pair was element 4 of group 6 (Fig.2b), corresponding to the best observing lateral resolution of $5.52\,\mu m$, which falls below the sensor’s pixel size of $5.94\,\mu m$. Furthermore, the evaluation metrics, i.e., the Fréchet inception distance (FID) and structural similarity index measure (SSIM), are 22.6855 and 0.8203, respectively, when comparing the reconstruction with the ground truth obtained by a bright-field microscope. Notably, using a higher-resolution sensor is a straightforward approach to improve lateral resolution.
We hypothesize that the diversity and quantity of samples in the training dataset also influence the final resolvable resolution. In the 1951 USAF resolution test, the observed object is only one test chart (see Methods), which will likely cause the over-fitting problem (see the failure case of singular sample in Fig.S7 in Supplementary Information). To address this, we conducted another experiment with $10\,\mu m$ diameter beads, increasing data diversity and quantity. The reconstructions clearly recover and resolve the bead shapes from different regions (see Fig.3a). Unlike traditional LFI confined to a thin focal plane, our results demonstrate a higher depth of field. This allows us to distinguish objects at various depths within the water flow (see Methods). We visualize the cross-sectional profiles of three beads at different depths in Fig.3a, along with their depth information encoded by pixel intensity. This capability for depth recovery paves the way for potential 3D high-throughput imaging using GenLFI.
Figure 3: Depth of field tests. a, $10\,\mu m$ diameter beads imaging. The beads, which were black and opaque, were reconstructed by LensGAN using their features from small dirt and transparent droplets. We select 3 beads at different depths and visualize them by cross-sectional profiles. In this particular case, the 3D depth map can be visualized based on pixel intensities from the reconstruction. b, Results of handwritten digits slides stack. Left shows the 4x objective focuses on “1” and is gradually out of focus from “2” to “4”. Right is the reconstructed result; all numbers are focused.
Multi-focus: Unlike conventional light microscopes requiring manual focus adjustments, GenLFI excels at reconstructing in-focus images from holograms containing objects at different depths. This unique capability is attributed to its virtual depth of field, offering significant flexibility for multi-focal imaging. To showcase this advantage, we conducted experiments with two stacks of 4 glass slides ($1\,mm$ thick), each marked with a written digit (font-weight $0.3\,mm$). In traditional bright-field setups, focusing is limited to a single depth plane, resulting in only one clear digit at a time (see Fig.3b). With proper data guidance (see Methods), LensGAN can refocus and resolve all digits across different slides in a single reconstruction, achieving a superior depth of focus of 3 mm (Fig.3b). This indicates our system’s versatility and underscores its potential in scenarios where simultaneous multi-depth clarity is crucial.
Inference time: Once deployed, the trained model relies only on a single network (i.e., the generator $G$ built by a 9-layer ResNet [47]) for reconstruction, making it the fastest among existing hologram reconstruction algorithms. Previous studies [17, 30] have suggested that DL can expedite the reconstruction process through an end-to-end approach. Unlike many DL-based reconstruction algorithms that rely on the multi-height phase recovery (MHPR) approach, LensGAN distinguishes itself by eliminating the necessity for multiple ($I$ number) hologram inputs for reconstruction. Notably, efficient model compression techniques for GANs [48] could further reduce the computational cost by 46×46$\times$46 ×, but we maintain the original model size to avoid resolution degradation. We evaluated LensGAN’s performance on our edge device (see MethodsMethods) using hologram images with $512\times 512$ pixels ($px$). LensGAN was compared against the self-supervised model GedankenNet [17], the supervised model U-Net [37], and the iterative algorithm MHPR [38], as outlined in Table 1. LensGAN achieved an equivalent inference time of $0.0031\,{s\cdot mm}^{-2}$ for a $1\,mm^{2}$ area of our microfluidic chip, showing its exceptional speed and efficiency in hologram reconstruction.
Table 1: Holographic image inference time (for $I$ input hologram(s), $512\times 512\,px$)
Method LensGAN ($I$=1) GedankenNet ($I$=2) U-Net ($I$=2) MHPR ($I$=2)
[…]
#Parameters 33.8M 39.4M 31.0M -
00footnotetext: M: million. Evaluations were performed on a single PC with an i9-13900K CPU and an RTX 3090 GPU.
Large FOV imaging of microfluidics
Figure 4: Reconstruction of a microfluidic chip ($30\times 5\,mm$). a, ROI in the hologram, i.e., the functional region containing dynamic droplets. b, Synthesized reconstruction result of the ROI (also see Supplementary video 1). We also analyzed the cross-sectional profiles at the center of 3 patches, labeled $i$, $o$, and $\hat{o}$, as denoted by the white dash line at $i$.
Microfluidic platforms have transformed biomedical research by enabling high-throughput experiments with biological samples down to the scale of single cells or molecules [49]. Despite their increasing adoption, a critical challenge remains. Existing high-resolution imaging techniques cannot simultaneously monitor multiple ROIs on a single chip. For instance, imaging the formation of droplets in droplet-based microfluidics often impedes simultaneous observation of downstream operations, such as droplet sorting or merging, occurring in a different ROI [49].
Our GenLFI approach overcomes this limitation. We demonstrate its capabilities in a microfluidic experiment, generating and dynamically imaging droplets across various chip sections concurrently. Data collection for the training dataset of this microfluidic experiment (see Methods) took approximately 20 minutes. Unlike most LFI systems, this procedure is automated, requiring no manual annotation. We acquired two-domain training data using our LFI system and a conventional bright-field microscope, labeled as $i$ and $o$, respectively (see Fig.1b). These two-domain images, i.e., $i$ and $o$, constituted the ‘GenLFI-Microfluidics’ dataset for LensGAN training and testing. Reconstruction results (see Fig.4) demonstrate that the trained LensGAN can generate high-resolution ($1792\times 768\,px$) bright-field images ($\hat{o}$) across the entire ROI (i.e., spanning a full microfluidic chip).
The results show that our approach not only preserves the global structure from the original hologram to the generated image but also ensures consistent transverse x-coordinate positioning of features, as evident in the re-focusing process shown in Fig.4b. Furthermore, the synthesized large FOV image (Fig.4b) reveals that although the trained LensGAN model infers each patch independently (see Methods), the microfluidic channels are aligned consistently across patches. This suggests the robustness of our approach in handling large-scale reconstruction tasks. This capability can facilitate monitoring and visual-based control for intricate microfluidic systems.
Large FOV imaging of spheroids
Figure 5: Reconstruction of cell spheroids. a, MeT-5A NF2KO spheroid images acquired using different imaging systems: bright-field microscopy with a 10x objective lens (left), digital holographic microscopy (DHM) with a 10x objective lens (center), and our GenLFI platform (right). b, Full-FOV reconstruction by LensGAN on an unseen well, post-processed by global panorama stitching algorithm (see Supplementary Information 6) and local image-stitching [50]. The structures are readily resolvable in the penetrated light parts at the edges of spheroids.
Cell spheroids, as straightforward yet extensively utilized multi-cellular 3D models, emerge from the natural inclination of most adherent cells to cluster together [51]. The imaging of spheroids poses significant challenges due to their considerable thickness, which can reach up to hundreds of microns, and their high density. As noted previously, conventional holographic imaging systems, including LFI systems, involve complex optical fields, resulting in contaminated holograms with overlapping scattered patterns and shadows (see Fig.1a). When utilizing digital holographic microscopy (DHM) with a 10x objective lens and a $635\,nm$ wavelength laser light source, only limited structural and intensity information can be retrieved (see Fig.5a).
With the addition of a light contrast-enhancing polarization filter, a bright-field microscope equipped with a 10x objective lens can discern a partial sectional structure at the periphery by adjusting the focus of the objective lens (Fig.5a). A comparable light penetration capability can also be achieved using our LFI platform by employing a red laser light source with a wavelength of $635\,nm$. Therefore, we integrated inverted bright-field microscopy to acquire data guidance images $o$, facilitating holographic reconstruction of similar quality but across an ultra-large FOV.
Fig.5b shows that our approach can reconstruct the spheroid edge structure and shapes from the measured holograms in a large FOV of an engineered cancer model. Notably, spheroids exhibit lateral structures that might replicate some of the more complex cellular environments, thereby mimicking more relevant scenarios while using high throughput platforms. Leveraging the advantages of the large FOV, our approach proves effective for monitoring the multiparametric spheroid growth patterns, which will likely allow for insights into biomechanical features, cell-cell interactions, and readily quantifiable multiparametric features of spheroid outgrowth assays in a scalable and cost-effective manner.
Discussion
GenLFI, a generative deep learning framework for LFI, eliminates the need for traditional forward imaging models by introducing data-driven guidance through unsupervised learning. This framework enables imaging in dynamic, noisy optical fields, previously not achievable with LFI. GenLFI achieves an over 20 times larger FOV than current LFI systems, enabling real-time monitoring for in vivo (e.g., zebrafish, see Fig.S1 in Supplementary Information) or in vitro (e.g., cell-in-droplets, 3D cell models) studies. Images are reconstructed using a DL network, significantly faster than traditional iterative techniques [22, 23, 52, 53, 38], while achieving sub-pixel resolution without requiring adjustments to the light source. These advancements demonstrate the potential of leveraging generative AI to overcome optical limitations.
The dynamic nature of our imaging scenarios makes it impractical to obtain paired ground truth and raw holograms for training. Therefore, GenLFI leverages unsupervised learning algorithms, inspired by those designed for image-to-image (I2I) transformation tasks in computer vision. This approach prioritizes hardware-agnostic and reproducible workflows with simple setups. While the current implementation utilizes a telescope cold camera with a resolution of $5.9\,\mu m/px$, GenFLI is compatible with higher-resolution sensors and multi-angle illumination for potential 3D LFI capabilities.
Our investigation into the physical correlation between holograms and bright-field images achieved leading results based on FID and SSIM metrics. However, challenges remain with such I2I-based approaches, primarily concerning data quality, quantity, and diversity. These factors directly impact reconstruction results, as demonstrated by artifacts observed in reconstructions from limited datasets (see the results from single-sample, paired datasets Fig.2b and Fig.S7 in Supplementary Information). Additionally, training GANs can be challenging due to potential issues like model collapse and non-convergence. While exploring diffusion models [54, 55, 56, 57] as an alternative shows promise, their limitations in handling unpaired and unaligned image translation tasks require further investigation. Furthermore, existing evaluation metrics like FID and KID primarily assess image-level distributions and lack the ability to evaluate the alignments between objects across different image patches. Developing new metrics accurately reflecting alignment between different images captured from the same objects is crucial for future advancements.
The combination of GenLFI and large vision models (LVM) [58, 59] holds promise for achieving a more generalized reconstruction capability for various samples and light field conditions. In addition, leveraging processors with advanced and specified architectures [60, 61, 62] could enable real-time image reconstruction with these large models.
Despite limitations due to optical conditions, GenLFI’s high-throughput capability opens doors for various biomedical applications. GenLFI offers a large FOV, enabling rapid analysis of large biological samples with minimal equipment. This capability is particularly valuable in resource-limited settings lacking traditional histopathology services. By incorporating digital staining through generative deep learning, GenLFI can bring advanced imaging techniques to such settings. Furthermore, GenLFI’s cost-effective and compact setup facilitates routine monitoring of cell culture and spheroid growth directly within incubators, reducing phototoxicity while enabling continuous monitoring of dynamic cellular processes across a large FOV. The testing of drug libraries based on dynamic processes, utilizing single-cell/cell models or droplet-based microfluidic systems, can now be obtained simultaneously from microplate-like platforms, paving the way for developing next-generation, autonomous, high-throughput drug screening systems.
Methods
Optical setup
GenLFI utilizes a versatile optical setup for capturing holograms. The optical setup comprises a light source, a CMOS sensor, and a test bench to stabilize their positions. We employed two types of light sources, including a blue LED light source for enhanced resolution (used for microfluidics, test chart, beads) and a red laser light source for deeper penetration depth (used for spheroids). For LED illumination, the light source is Thorlabs M420L3 (wavelength $420\,nm$) blue LED, with power supply (15V 1A) and T-Cube LED Driver (Thorlabs LEDD1B). For laser illumination, we used an optical fibre-coupled red laser source (Thorlabs S1FC635, wavelength $635\,nm$). The mechanical setup consists of a cage plate (Thorlabs CP33/M), a mounting bracket (ThorLabs 973/579-7227), a $25\,mm$ linear stage (Thorlabs XR25C) with an end-mounted micrometer, a post holder (Thorlabs LPH200), a post (Thorlabs P350/M), and an adjustable platform (Thorlabs C1519). The telescope camera (ZWO ASI 2400MC Pro) features a full-frame Sony IMX410 CMOS sensor, $36\times 24\,mm$ in size, with a $5.94\,\mu m/px$ resolution. Mounted on an experimental bracket (FAMA MV test stand), the camera aligns with the light source and the imaging object along a single axis. The distance from the light source to the object’s bottom plane is $430\,mm$, and from the object’s bottom plane to the CMOS sensor is $12.5\,mm$.
LensGAN architecture
LensGAN, a core component of GenLFI, reconstructs high-resolution images from captured holograms. LensGAN generally considers the measured and reconstructed holograms ($i$ and $\hat{i}$) and the measured and generated bright-field images ($o$ and $\hat{o}$) into two domains, $X$ and $Y$, respectively. It leverages a GAN framework with a cycle-consistent training strategy to learn the complex mapping between holograms (domain $X$) and corresponding bright-field images (domain $Y$). The architecture of LensGAN comprises 6 networks: generators of reconstruction and reverse mappings $G:X\rightarrow Y$ and $F:Y\rightarrow X$, selectors $\beta_{X}$ and $\beta_{Y}$, and discriminators $D_{X}$ and $D_{Y}$. During training, all 6 networks function following the training loop as shown in Fig.1a. In the testing phase, only the trained generator $G$ is used for reconstruction. The outputs of LensGAN are reconstructed bright-field images $\hat{o}=G(i)$ and generated holograms $\hat{i}=F(o)$.
Generators $G$ and $F$ handle the fundamental image transformation task, while the remaining 4 networks contribute to the training process. $G$ and $F$, each composed of 9 residual blocks (ResNet [47]), feature an image encoder for extracting image characteristics and a decoder for generating images. The selector network $\beta$ [63] is responsible for pairing images based on feature alignments and is constructed with an adaptive average pooling layer, 4 convolutional networks, and one fully connected network (FCN) with a Softmax layer. The discriminator $D$ is designed to differentiate between images, emphasizing structural information and intricate details within the complex amplitude. Recent advancements in Transformers [64] have demonstrated the equivalence of multi-layer and one-layer attention mechanisms [65]. By leveraging this finding, we designed $D$ with a Markovian architecture, merging a one-layer attention module with 5 downsampling convolutional layers to augment amplitude details while maintaining minimal computational overhead (see Supplementary Information 4, 5).
Combining these networks, the training loop adheres to the cycle structure [66], guided by several loss functions, depicted in Fig.1b and Fig.S2 in Supplementary Information. To augment phase structural feature information during the training, we devised a phase contrast module positioned to constrain the physical correlation between $X$ and $Y$ domains. The observed samples in LFI are generally light-permeable, resulting in duotone holograms captured under a single-wavelength light source. In such scenarios, phase contrast [6] and dark-field microscopes are commonly employed, where a phase ring and other lenses are combined to create a phase contrast. Inspired by this technique, our approach is to develop a module that learns the phases of the two domains contrastively, thereby mimicking a similar mechanism and enhancing the physical correlations between the object’s hologram and bright-field images. Technically, we create the phase-contrast loss function $\mathcal{L}_{\mathrm{PC}}$ by a multi-scale structural similarity index measure (MS-SSIM) [67] for the amplitude spectra:
$\min_{G}\mathcal{L}_{\mathrm{PC}}(X,Y)=\mathbb{E}_{i\sim P_{X}}\left[1-\operatorname{MS-SSIM}(|\mathcal{FT}(i)|,|\mathcal{FT}(\hat{o})|)\right],$ (3)
where $|\mathcal{FT}(\cdot)|$ denotes the magnitudes of the individual frequency components obtained from the Fourier transform ($\mathcal{FT}$) of an image, and $P_{X}$ is the probability distribution for $i$. $\operatorname{MS-SSIM}$ is defined as:
$\operatorname{MS-SSIM}(x,y)=\left[l_{M}(x,y)\right]^{\gamma_{M}}\cdot\prod_{j=1}^{M}\left[c_{j}(x,y)\right]^{\delta_{j}}\left[s_{j}(x,y)\right]^{\epsilon_{j}},$ (4)
with the terms
$l(x,y)=\frac{2\mu_{x}\mu_{y}+C_{1}}{\mu_{x}^{2}+\mu_{y}^{2}+C_{1}},$ (5)
$c(x,y)=\frac{2\sigma_{x}\sigma_{y}+C_{2}}{\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2}},$
$s(x,y)=\frac{\sigma_{xy}+C_{3}}{\sigma_{x}\sigma_{y}+C_{3}},$
where $x$ and $y$ and two input image patches, $M=5$ denotes the highest scale number set for the system to iteratively apply a low-pass filter between $x$ and $y$. $\mu$ and $\sigma^{2}$ stand for mean value and variance, respectively, which are implemented by 2D convolutions. $l(x,y)$, $c(x,y)$, and $s(x,y)$ represent luminance, contrast, and structure comparison measures between $x$ and $y$, with control weights $\delta_{1}=\epsilon_{1}=0.0448$, $\delta_{2}=\epsilon_{2}=0.2856$, $\delta_{3}=\epsilon_{3}=0.3001$, $\delta_{4}=\epsilon_{4}=0.2363$, and $\gamma_{5}=\delta_{5}=\epsilon_{5}=0.1333$. LensGAN enforces all inputs as 8-bit images with a pixel value range of 0 to 255; the constants are set as $C_{1}=(0.01\times 255)^{2}$, $C_{2}=(0.03\times 255)^{2}$, and $C_{3}=C_{2}/2$.
Note that in (3), we added $\mathcal{L}_{\mathrm{PC}}$ only for the mapping $G:X\rightarrow Y$, the transformation process of holographic reconstruction, with a control weight $\tau_{\mathrm{PC}}$. Similarly, only for $G$, selector $\beta_{X}$ is regularized by the effective sample size loss [63] for $\beta_{X}$:
$\min_{\beta_{X}}\mathcal{L}_{\mathrm{ESS}}(X)=\|\beta_{X}\|_{2},$ (6)
where $\|\cdot\|_{2}$ denotes the L2 norm.
For the whole cycle (reconstruction mapping $G:X\rightarrow Y$ and reverse mapping $F:Y\rightarrow X$), there is a GAN loss function $\mathcal{L}_{\mathrm{GAN}}$ to constrain $G$, $F$ and $D$, with two consistency loss functions $\mathcal{L}_{\mathrm{cyc}}$ and $\mathcal{L}_{\mathrm{idt}}$ to further regularize $G$ and $F$ [68, 66, 63]. They are reweighted least squares adversarial loss
$\min_{G,\beta_{X}}\max_{D_{Y}}$ $\mathcal{L}_{\mathrm{GAN}}(X,Y)$ (7)
$=$ $\mathbb{E}_{i\sim P_{X}}\beta_{X}(i)\left[\left(1-D_{Y}(\hat{o}\right)^{2}\right]+$
$\mathbb{E}_{o\sim P_{Y}}\beta_{Y}(o)\left[D_{Y}(o)^{2}\right],$
reweighted cycle consistency loss
$\min_{G,F}\mathcal{L}_{\mathrm{cyc}}(X,Y)=\mathbb{E}_{i\sim P_{X}}\beta_{X}(i)\|i-F(\hat{o})\|_{1}$ (8)
$+\mathbb{E}_{o\sim P_{Y}}\beta_{Y}(o)\|o-G(\hat{i})\|_{1},$
and reweighted identity loss
$\min_{G,F}\mathcal{L}_{\mathrm{idt}}(X,Y)=\mathbb{E}_{i\sim P_{X}}\beta_{X}(i)\|i-F(i)\|_{1}$ (9)
$+\mathbb{E}_{o\sim P_{Y}}\beta_{Y}(o)\|o-G(o)\|_{1},$
where $P_{X}$ and $P_{Y}$ are the probability distributions for $i$ and $o$, respectively, and $\|\cdot\|_{1}$ denotes the L1 norm.
The LensGAN’s full objective of loss functions is combined with:
$\begin{gathered}\mathcal{L}_{main}=\min_{G,F}\max_{D_{X},D_{Y}}\mathcal{L}_{\mathrm{GAN}}(X,Y)+\mathcal{L}_{\mathrm{GAN}}(Y,X)\\ \quad+\tau_{\mathrm{cyc}}\mathcal{L}_{\mathrm{cyc}}(X,Y)+\tau_{\mathrm{PC}}\mathcal{L}_{\mathrm{PC}}(X,Y)+\tau_{\mathrm{idt}}\mathcal{L}_{\mathrm{idt}}(X,Y),\\ \mathcal{L}_{\beta_{X}}=\min_{\beta_{X}}\mathcal{L}_{\mathrm{GAN}}(X,Y)+\tau_{\mathrm{ESS}}\mathcal{L}_{\mathrm{ESS}}(Y),\\ \mathcal{L}_{\beta_{Y}}=\min_{\beta_{Y}}\mathcal{L}_{\mathrm{GAN}}(Y,X)+\tau_{\mathrm{ESS}}\mathcal{L}_{\mathrm{ESS}}(Y),\end{gathered}$ (10)
where $\mathcal{L}_{main}$ is used to train networks $G$, $F$, and $D$, $\mathcal{L}_{\beta_{X}}$ and $\mathcal{L}_{\beta_{Y}}$ are used to train $\beta_{X}$ and $\beta_{Y}$, respectively; the weights $\tau_{\mathrm{cyc}}$, $\tau_{\mathrm{PC}}$, $\tau_{\mathrm{idt}}$, and $\tau_{\mathrm{ESS}}$ control the relative importance of different losses and can be adjusted as hyperparameters.
Training, testing and post-processing
For all experiments, we set hyperparameters $\tau_{\mathrm{cyc}}=10$, $\tau_{\mathrm{PC}}=8$, and $\tau_{\mathrm{ESS}}=\tau_{\mathrm{idt}}=1$. We use Adam solver [69] to initiate training of the networks from scratch, maintaining a consistent learning rate of 0.0001 for the initial 50 epochs. Subsequently, we linearly decrease the learning rate to zero over the following 50 epochs. Each epoch processed 10,000 images (we set the input dimension as $256\times 256\,px$) with a batch size 20. The training was on a single Nvidia RTX A100 GPU; for dataset ‘GenLFI-Microfluidics’, it took 29 hours. All tests were conducted on a workstation (CPU Intel i9-13900K, GPU Nvidia RTX3090). We utilized FID for quantification and comparative evaluation. Following fast reconstruction, we performed post-processing for large sample results using image stitching algorithms. As illustrated in Fig.5b, we employed local object stitching [50] and global panorama stitching (see Supplementary Information 6) separately.
Microfluidics imaging
We fabricated a microfluidic chip with a flow-focusing geometry and conducted water-in-oil droplet generation experiments to observe these dynamic samples. The microfluidic channel width and height are $200\,\mu m$ and $35\,\mu m$, respectively. We used piezoelectric disc pumps (TTP Ventus HP series) to control the microfluidic flows and droplet production rates (ranging from 5 to 15 $Hz$). We conducted the imaging experiments under a bright-field microscope (Cerna modular microscopy with Thorlabs CS135MU camera) and our LFI platform, respectively, and acquired image data $i$ and $o$ for network training.
Using the GenLFI setup, we recorded 500 hologram frames ($i$) in total (.fit format, resolution $6072\times 4042\,px$, acquired via the ASI Studio software) focusing on the ROI of the microfluidic chip with dynamic droplets (see Fig.4). Then, under the same experimental conditions, we repeated the experiment to acquire real-world bright field images ($o$) as the data guidance using a bright-field microscope. Due to the small FOV of the microscope, we scanned the same ROI and recorded a 5-minute video containing 7000 frames of bright-field images.
We cropped the hologram images to patches with a size of $166\times 147\,px$. Similarly, bright-field images were cropped to patches with $1024\times 1080\,px$. The final dataset ‘GenLFI-Microfluidics’ for this task comprises 121,886 training and 100 test cropped images from LFI hologram ROIs, and 57,193 training and 100 test cropped images from bright-field images.
Cell spheroids imaging
Human pleural mesothelial cells MeT-5A NF2KO were cultured as 3D spheroids by seeding approximately 30,000 cells into each well of a 6-well ultra-low attachment culture plate (#3471; Corning, Thermo Scientific); details see [70]. The dataset ‘GenLFI-Spheroids’ construction and network training for this task follows a similar routine of Microfluidics imaging. We used a red laser (see Optical setup) as the light source. The bright-field microscope for recording is Leica DMIRE2 with a 10x objective lens, coupling with a polarization filter to enhance the light contrast for spheroid structure observation. The acquired raw data contains 132 frames of holograms and 26,581 frames of bright-field microscopic images. We constructed a dataset comprising 17,418 training and 100 test images (captured from an unseen well) for LFI holograms (after Bicubic interpolation super-resolution, $808\times 808\,px$), and 23,738 training and 100 test images for bright-field images (original size, $1280\times 1024\,px$).
USAF 1951 test chart imaging
We employed a USAF 1951 test chart (Thorlabs R3L3S1P) to evaluate the resolution limits of our LFI system. Given that the thickness of the test bench’s glass slide, i.e., $1.5\,mm$, affects the light diffraction, we placed the printed side close to the microscope’s lens to minimize this effect when capturing images through standard bright-field microscopy. Then, we placed the sample in the same orientation to acquire holograms with GenLFI. For microscopy images, we selected a square ROI with the smallest pair of lines. For holograms, we cropped out the same ROI to a $220\times 220\,px$ image (see Fig.2b). We carefully aligned these two domain images on the same scale. To augment our dataset for this task, we introduced Gaussian noise to generate 10,000 variant images for each domain, of which 99 images with 1 original noise-free image as test images, and the other images as training images.
Beads imaging
We used $10\,\mu m$ diameter beads solution (Invitrogen CML latex, $4\%$ w/v, H2392891) diluted with water on a microscope slide for imaging. The water flow on the slide is approximately $0.1\,mm$ thick; therefore, the $10\,\mu m$ beads float or sink across the flow. We recorded 23 frames of hologram images ($6072\times 4042\,px$) under the LFI platform, and 1430 frames ($4096\times 2160\,px$) images under the bright-field microscope with a 4x objective. For both recorded holograms and bright-field images, we applied Gaussian filtering to remove shadows, then cropped and applied data augmentation with vertical and horizontal flips to scale up the data size. The dataset, ‘GenLFI-Beads’, consists of 96 test and 33,120 training images of hologram ($506\times 505\,px$); 96 test and 32,576 training images of bright-field images ($512\times 540\,px$).
Multi-focus imaging
In evaluating multi-focus imaging capabilities, we stacked four glass slides, each $1\,mm$ thick, upon which handwritten numbers (1,2,3,4) had been inscribed. Another slack is marked with 5,6,7,8 at different heights as the unseen object for the test. These numbers serve as markers at different distances from the CMOS sensor’s imaging plane. When examined under bright-field microscopy, focusing on any one number caused the others to appear progressively more out of focus. We captured focused images of each number separately to establish clear ground truth data for training the LensGAN. The last dataset, ‘GenLFI-Multifocus’, consists of 43 test and 100 training images of hologram (1072×(669 or $1072\times(669\text{ or }1338)\,px$); and 27 test and 395 training images of bright-field images ($4096\times 2160\,px$). During training, hologram images were guided solely by these focused images.
    Supplementary Information
    1 Zebrafish imaging
    We applied GenLFI to in vivo samples of zebrafish. Zebrafish were cultured to 3 days old, anesthetized by MS222, and placed under petri dishes to be observed. Due to the structure of zebrafish, the measured hologram on our sensor (resolution $5.94\,\mu m/px$) can provide very limited information inside the fish (see Fig.S1). Yet, our system can still reconstruct the zebrafish’s basic structural information from raw holograms in a large FOV.
    Figure S1: a, Details of imaging for zebrafish; b, Synthetized full FOV images.
    2 Further discussion on complex optical field
    Here we further discuss the optical field with large axial size samples and mathematically prove the inadequacy of conventional model-based 3D lens-free holographic microscopy (LFHM) tomographic reconstruction approaches. To this end, we analyze the current 3D LFHM algorithm based on scalar diffraction theory for 3D objects.
    As shown in Fig.1 in the main text, the solution to the light propagation problem is described as:
    $u(\boldsymbol{r})=u^{inc}(\boldsymbol{r})+u^{sct}(\boldsymbol{r})$ (11)
    where $u^{inc}(\boldsymbol{r})$ represents the incident light and $u^{sct}(\boldsymbol{r})$ represents the scattered light. The Lippmann-Schwinger equation (LSE) depicts the relationship between the 3D scattering potential of the sample and the resulting scalar field at region $V$:
[…]
    where $G(\boldsymbol{r})=\frac{1}{4\pi R}\exp(jk_{b}R)$, $k_{b}=k_{0}n_{b}$ is the wavenumber of the light in the background medium, and $R$ denotes the length of the Euclidean vector $\boldsymbol{r}$. Discretized and denoted by vectors, the resulting scalar field $\boldsymbol{\hat{u}}$:
    $\widehat{\boldsymbol{u}}=\widehat{\boldsymbol{u}}^{inc}+\widehat{\boldsymbol{G}}\boldsymbol{\operatorname{diag}(f)}\boldsymbol{u}^{inc}=\mathcal{L}(\boldsymbol{f})$ (13)
    where $\mathcal{L}(\cdot)$ is the light propagator, $\boldsymbol{\operatorname{diag}(f)}\in\mathbb{C}^{N\times N}$ denotes a diagonal matrix whose main diagonal entries consist of elements from the vector $\boldsymbol{f}$. With multiple light sources, the final forward model of the 3D LFHM can be presented by the intensity of the optical field $\widehat{\boldsymbol{H}}$ under $p^{th}$ illumination:
    $\widehat{\boldsymbol{H}}=\left|\widehat{\boldsymbol{u}}^{inc}+\widehat{\boldsymbol{G}}\boldsymbol{\operatorname{diag}(f)}\boldsymbol{u}^{inc}\right|^{2}=\left|\mathcal{L}_{p}(\boldsymbol{f})\right|^{2}$ (14)
    The reconstruction algorithm can be formulated as the optimization problem:
    $\boldsymbol{f}^{*}=\arg\min_{f\in\mathbb{C}^{N}}\{\Sigma_{p=1}^{P}d(\left|\mathcal{L}_{p}(\boldsymbol{f})\right|^{2},\widehat{\boldsymbol{H}}_{p})\},$ (15)
    where $\mathcal{L}_{p}(\cdot)$ denotes the light propagator, $\boldsymbol{f}^{*}$ denotes the amplitude of the light, $\widehat{\boldsymbol{H}}_{p}$ represents the measured light intensity under the $p^{th}$ illumination.
    In the microfluidics experiment, the mediums consist of air, PDMS, water, oil, and glass. In this condition, different components of the electromagnetic field may make the vector diffraction theory of $u(\boldsymbol{r})$ in (11) ineffective. Furthermore, the model-based algorithm’s approximation of the hologram in the 2D space is affected by the thickness of the sample. For example, the Born approximation for (13) is valid only if $\boldsymbol{u}^{sct}\ll\boldsymbol{u}^{inc}$. This corresponds to $l(n_{s}-n_{b})\ll\lambda$ , where $l$ is the thickness, $(n_{s}-n_{b})$ describes the medium’s refractive index variation, and $\lambda$ is the wavelength of the incident light. Consider a PDMS chip that has water inside, with a thickness of $5000\,\mu m$; there are $l(n_{s}-n_{b})$ = $5000\times(1.40-1.33)=350\,\mu m$, while $\lambda=0.42\,\mu m$. It is obvious that the model is not applicable to the condition of Born approximation for large samples with strong scattering.
    To provide a more accurate description of the physical process, some non-linear models have been proposed [30]. These iterative methods involve either multiple iterations by variations or dividing the sample into a series of thin layers. However, these methods are impractical to apply in real experiments and are restricted to sample-specific tools.
    In summary, real-world scattered fields are complex and difficult to model. The existing forward model-based LFI systems are limited to small and thin samples.
    3 Full architecture of LensGAN
    Figure S2: The full architecture of LensGAN.
    4 Discriminator design
    Figure S3: Architecture of the 5-convolutional-layer discriminator. f: filter, k: kernel size, s: stride, p: padding. The self-attention layer calculates the self-attention scores from 3 vectors: query, key, and value. Then it uses the scores to create an attended feature map that is added back to the input.
    The discriminator network is based on $70\times 70$ PatchGANs [71], which aims to classify whether $70\times 70$ overlapping image patches are real or fake. The input images under our hyperparameter setting are $256\times 256$, then downsampled by convolutional layers. The output is a one-channel prediction map to discriminate images. As shown in Fig.S3, the convolutional layers only process information in local neighborhoods. The sequential Markovian modeling of long-range dependencies in images may cause a limited receptive field. Therefore, a self-attention layer is added to solve this problem and make the training more robust.
    We implement this self-attention layer with the feature map used in SAGAN [72]. The input image from the previous convolutional layer contains features $x\in\mathbb{R}^{C\times N}$ that are transformed into two feature spaces $\mathnormal{f}$ and $\mathnormal{g}$. Here, $C$ represents the number of channels, and $N$ represents the number of feature locations. The attention function is calculated by the following equation where $f(x)=W_{f}x$ , $g(x)=W_{g}x$ :
    $\beta_{j,i}=\frac{\exp(s_{ij})}{\sum_{i=1}^{N}\exp(s_{ij})},\text{where }s_{ij}=f(x_{i}^{T})g(x_{j})$ (16)
    $\beta_{j,i}$ indicates the range of attention where the model computes mapping of the $i^{th}$ location of the $j^{th}$ feature regions. The output of the attention layer is $a=(a_{1},a_{2},...,a_{j},...,a_{N})\in\mathbb{R}^{C\times N}$ , where
    $a_{j}=v(\sum_{i=1}^{N}\beta_{j,i}h(x_{i})),h(x_{i}=W_{h}x_{i}),v(x_{i}=W_{v}x_{i})$ (17)
    $W_{g}\in\mathbb{R}^{\overline{C}\times C}$ , $W_{f}\in\mathbb{R}^{\overline{C}\times C}$ , $W_{h}\in\mathbb{R}^{\overline{C}\times C}$ , and $W_{v}\in\mathbb{R}^{C\times\overline{C}}$ are three learned weight metrics that are implemented as $1\times 1$ convolutions. Practically, for memory efficiency, the channel number of $\overline{C}$ is reduced to be $C/k$ , where $k$ is set as 8. By multiplying $a_{i}$ a learnable scale parameter $\gamma$ and adding back the input feature map $x_{i}$, the final scaled output is:
    $y_{i}=\gamma o_{i}+x_{i}{.}$ (18)
    5 Ablation study
    Figure S4: Visual comparison of results by different I2I algorithms. We selected 3 cases from the ’Hologram2Microscopy’ dataset and evaluated different algorithms with FID: classical iterative holographic reconstruction algorithm [22], our algorithm LensGAN, CycleGAN [66], our backbone IrwGAN [63], IrwGAN enhanced with our self-attention discriminator, and SOTA unpaired I2I algorithm GP-UNIT [73].
    We evaluated LensGAN on an unseen test set (images from full ROI frames) with Fréchet Inception Distance (FID) [74], as depicted in Fig.S4. FID is the current standard for assessing the quality of generative models, where a lower FID indicates a better correlation between the distributions of generated images and ground truth. In the ablation study, we replaced the discriminator in IrwGAN [63] with our own (enhanced with self-attention blocks, details in Discriminator design), denoted as ‘IrwGAN + Self-attention.’ We also evaluated the CycleGAN [66] and a SOTA unpaired I2I model GP-UNIT [73] on our dataset.
    Fig.S4 shows the results that compared to the IrwGAN [63] backbone, the new discriminator improves the FID score by 3.27, and together with the $\mathcal{L}_{\mathrm{PC}}$, the LensGAN improves the FID score by 10.41. The results in Fig.S4 demonstrate that our model achieved the lowest score and successfully addressed the main distortions and artifacts. $\mathcal{L}_{\mathrm{PC}}$ constrains the generated results effectively, preventing the line artifacts in IrwGAN (see cases 1 & 3), and our discriminator $D$ improves the reconstructed droplets’ illumination details (see case 2).
    6 Post-processing: Global panorama image stitching
    After processing the image blocks separately, the results will be stitched together to create the final panorama. This operation is divided into two stages. The first stage involves intensity correction and image assembly. The second stage aims to remove slowly varied artifacts.
    For the first stage, considering two image blocks to be stitched, the color difference usually occurs between them, causing a noticeable seam. We define one image as the reference image $I^{r}$ and the other image as the target image $I^{t}$. Pixel points at the seam line are denoted by $p_{i},i=1,2,...,n$. Then the color difference $\Delta C(p_{i})$ between $I^{r}$ and $I^{t}$ at the seam line is defined as:
    $\Delta C(p_{i})=V^{r}(p_{i})-V^{t}(p_{i}),~{}~{}i=1,2,...,n$ (19)
    where $V^{r}(p_{i})$ denotes the average pixel intensity at $p_{i}$ in the $I^{r}$ and $V^{t}(p_{i})$ represents the average pixel intensity at $p_{i}$ in the $I_{t}$. $V^{r}(p_{i})$ is calculated by averaging the pixel intensity at $p_{i}$ with its nearest three neighbors at the same row in the $I^{r}$. And the same approach is adopted to estimate $V^{t}(p_{i})$. For each pixel $p^{t}$ in the $I^{t}$, the intensity change $\Delta V$ is defined as [75]:
    $\Delta V(p^{t})=\frac{1}{\sum_{i=1}^{n}w_{i}}\sum_{i=1}^{n}w_{i}\Delta C(p_{i}),$ (20)
    where the weight $w_{i}$ is expressed as:
    $w_{i}=\mathrm{exp}\left(-\left(\frac{||V^{t}(p_{i})-I^{t}(p^{t})||^{2}}{\sigma_{1}^{2}}+\frac{||L(p_{i})-L(p^{t})||^{2}}{\sigma_{2}^{2}}\right)\right).$ (21)
    Then, the corrected intensity at $p^{t}$ is formulated as:
    $\overline{I^{t}(p^{t})}=I^{t}(p^{t})+\Delta V(p^{t}).$ (22)
    There exist slowly varied artifacts in the generated panorama at the first stage. In the second stage, we treat this type of artifact as the low-frequency component of the image. Therefore, we can use the Gaussian filter to estimate these artifacts. Suppose the initial panorama is denoted by $P$, the artifact image is estimated by:
    $A=P\ast\kappa,$ (23)
    where $\ast$ represents the convolution, and $\kappa$ is the Gaussian kernel. Therefore, the final panorama $\overline{P}$ is expressed as:
    $\overline{P}=P-A.$ (24)
    Figure S5: Stitching result for microfluidics. Up: fast reconstruction. Down: after panorama stitching.
    Figure S6: Stitching result for spheroids. Top left: input raw hologram; Bottom left: fast reconstruction from LensGAN; Top right: detected shadow; Bottom right: after global panorama stitching and shadow removing.
    7 Microfluidic chip fabrication
    We fabricated a microfluidic chip with a flow-focusing geometry for droplet production on PDMS bonded to a glass slide, and we conducted droplet generation experiments to observe these dynamic samples across a microfluidic system encompassing the droplet production area, a serpentine channel for droplet incubation and an expansion area for enabling droplet contact and merging. To fabricate the mold from which the microfluidic chip is made, photolithography is used. A $35\,\mu m$ thick layer of photoresist (SU-8 2015) is spin-coated, pre-baked, exposed, post-baked, and developed.
    After the silicone wafer mold containing the design is created, a PDMS base and curing agent (Sylgard 184) are combined in a 10:1 ratio in a plastic cup and stirred for several minutes using a spatula. The plastic cup is placed in a vacuum desiccator to extract trapped air in the PDMS/curing agent mix. After approximately 20 minutes the mix was poured over the earlier created master, after which it was placed in the vacuum desiccator until no bubbles were visible anymore within the PDMS/curing agent mix. After this, the mix was placed in an oven at $65^{\circ}$ for 3 hours to cure the PDMS.
    After the PDMS was cured, the PDMS was removed from the master, and separate devices were cut from this removed PDMS slab. Inlet and outlet holes were punched using a $1.2\,mm$ biopsy punch. Using scotch tape, dust was removed from both the glass slide and the PDMS devices, after which both were exposed to oxygen plasma in a plasma cleaner (Harrick PDC-002). The PDMS-coated glass slides and PDMS devices were then bonded by carefully connecting them.
    Finally, the PDMS chip is silanized using a 5 wsilane and HFE-7500 mixture.
    8 Spheroid genotypes
    The genotypes included are the wild type (WT) as well as two MeT5A clones with knockout (KO) of the NF2 gene – NF2 KO #1 and NF2 KO #2 (where #1 and #2 clones were generated using CRISPR constructs targeting different regions the gene). The NF2 gene has been found to have loss of function mutations in several cancers, including mesothelioma, and here these NF2 KO cell lines were generated as a cellular model system for mesothelioma.
    9 Failure case
    Figure S7: Failure case of singular sample: reconstruction for an organ-on-chip. We used a single raw hologram and BF microscopy image for training. Due to the low data scale, although the reconstruction achieved a high SSIM score, some information are lost in the reconstruction.